OpenStudy (anonymous):

help!! find the value of j if j=2a squared(b-3)squared ,a=3 and b=-1 answers please.

OpenStudy (kc_kennylau):

Do you mean \(j=2a^2(b-3)^2\)?

OpenStudy (kc_kennylau):

Just plug in the values inside

OpenStudy (kc_kennylau):

The first step would be \(j=2a^2(b−3)^2=2(3)^2(-1−3)^2\)

OpenStudy (kc_kennylau):

Do the calculations yourself

OpenStudy (anonymous):

need answers...i tried still but i couldn't...

OpenStudy (praxer):

it is 288 as, 2 X 3^2 X (-4)^2 = 2 X 9 X 16 =18 X 16 =288

OpenStudy (praxer):

@Raince did you get the steps

OpenStudy (anonymous):

yerp! thanks praxer :D

OpenStudy (praxer):

welcome :)